Its Telling me to solve this equation. i dont understand how to work this problem out i know i have to get a on one side to find out what it represents but i tried and dont get no where .
(5/4)(2a-10)=(-3/4a)

Answer by Alan3354(69443)

(5/4)(2a-10)=(-3/4a)
Maybe you mean
(5/4)(2a-10)=(-3/4)a
Multiply by 4
5(2a-10) = -3a
Expand
10a - 50 = -3a
Add 50
10a = -3a + 50
Add 3a
13a = 50
Divide by 13
a = 50/13